Consider the following hypothetical nuclear reacti

Please explain!! and answer all parts!!! I cant get anybody to completely answer the question!!
Answer #1

the above reaction is alpoha decay so

195 Pt78 -------------> 191W76 + 4 He 2

so X is Tungsten

A is 191

Z is 76

Q = M( 195 Pt78 ) - M(191W76)- M(4 He 2) c^2

= 194.96479 u - 190.96094 u -4.002603 u) c^2 ( 931.5 MeV/c^2 u)

=1.161 MeV

(c)

speed of the alpha particle

Q = 1/2 mv^2

v = root 2Q/ m = root 2( 1.161 MeV( 1.6* 10 ^-13 J/ MeV/6.64424* 10 ^-27 kg

=7.47 * 10 ^6 m/s

(d)

momnetum of alpha particle is

P = mv = 6.64424* 10 ^-27 kg(7.47 * 10 ^6 m/s) = 4.96* 10 ^-20 kg m/s